# Leray-Schauder Mappings for Operator Learning

### v2

@@ 3. Algorithm
Theorem 2.1 gives a foundation for the algorithm discussed in this section, whose implementation is the main objective of the present article. Our deep learning algorithm is based on constructing a function input from the data available for initialization (e.g. an initial condition or boundary values, depending on the formulation of the problem), and (nonlinearly) project the input on a basis of functions modeled with neural networks. The projection maps are Leray-Schauder maps, defined through Equation (1). The maps $\mu_{i}$ are either fixed as in Equation (3) (for some choice of norm), or are learned as in Theorem 2.2 and Theorem 2.5. In fact, while our experiments show that the assignment in Equation (3) gives a functioning model, we have found that hyperparameter fine tuning is more complex in this setting, and simply learning the $\mu_{i}$ maps, which in turn means that we learn the Leray-Schauder maps as well, makes training much simpler.
The model consists of $n$ feed forward neural networks $g_{i}$, $i=1,\ldots,n$, where each $g_{i}:K\longrightarrow\mathbb{R}^{M}$, and a feed forward neural network $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$. Since we also learn the functions $\mu_{i}$ in Equation (3), we also have convolutional neural networks $\mu_{i}:C(K,\mathbb{R}^{M})\longrightarrow\mathbb{R}^{+}_{0}$. In practice, we use CNNs to implement the $\mu_{i}$ neural networks because they mimic an integration procedure which can be performed in higher dimensions. They take a discretized input function and produce a numerical value (which can be constrained to be nonnegative by choice of the final nonlinearity).
The neural networks $g_{i}$ have the role of projection basis, since they are the functions whose span gives the finite dimensional approximation space in the proof of Theorem 2.5.
@@
which is the same operator used in the approximation of Theorem 2.5, where $F^{i}$ has been named $\mu^{i}$ to mimic the definition in (3). Since $P_{n}(\mathbf{y})$ can be evaluated through the networks $g_{i}$, it is clear that $P_{n}$ takes a function as input, obtained via initialization data, and produces a function as output. This function $P_{n}(\mathbf{y})$ lives inside the span $\langle g_{1},\ldots,g_{n}\rangle$. In fact, $P_{n}(\mathbf{y})$ is a linear combination of $g_{i}$, with coefficients $q_{i}:=\frac{\mu^{i}(\mathbf{y})}{\sum_{j=1}^{n}\mu^{j}(\mathbf{y})}$. The vector of coefficients $\mathbf{q}=(q_{1},\ldots,q_{n})^{T}\in\mathbb{R}^{n}$ is therefore the input of the network $f$, following the proof of Theorem 2.5.
We compute $\mathbf{b}=f(\mathbf{q})$, and $\mathbf{b}$ is used to take a linear combination of the neural networks $g_{i}$, which we set $\psi=\sum_{i=1}^{n}b_{i}g_{i}$. We can then evaluate $\psi(\mathbf{x},t)$ and compute the loss $\|\psi(\mathbf{x},t)-\mathbf{z}(\mathbf{x},t)\|$, where $\mathbf{z}$ is the target data which we are predicting.
This is summarized in Algorithm 1. We will refer to this method as the Leray-Schauder Neural Operator for brevity.
@@
1:Neural networks $\{g_{i}\}_{i=1}^{n}$ and $f_{n}$
2:Approximation $\phi_{n}^{-1}f_{n}\phi_{n}\tilde{P}_{n}$ of operator $T$
@@
6:Apply $f_{n}$ on $\mathbf{q}$: $f(\mathbf{q})=\mathbf{b}$
7:Output function, using $\phi^{-1}_{n}$, is $\psi(\mathbf{x},t)=\sum_{i=1}^{n}b_{i}g_{i}(\mathbf{x},t)$
@@ 4. Experiments
Leray-Schauder $0.0011\pm 0.0005$ $0.0011\pm 0.0005$ $0.0017\pm 0.0012$ $0.0017\pm 0.0011$
ANIE $0.0014\pm 0.0002$ $0.0015\pm 0.0003$ $0.0016\pm 0.0016$ $0.0017\pm 0.0018$
Spectral NIE $0.0022\pm 0.0016$ $0.0028\pm 0.0032$
FNO1D (init 5) $0.0292\pm 0.0285$ $0.0994\pm 0.1207$
@@
[…]
@@
In the experiments, we find that the model is comparable with the state of the art model (ANIE, [ANIE]) on both datasets. For the IE spirals, we have inlcluded two experiments, one on prediction of the original dynamics, and another one on an interpolation task where the model has access to half the points during training and predicts all of them during evaluation. For the Burgers’ dataset we have experimented on two spatial resolutions $s=256$ and $s=512$. An example of ground truth dynamics for the Burgers’ equation is found in Figure 1, while the corresponding model’s prediction is in Figure 2. The vertical direction indicates space and the horizontal direction indicates time.
@@
We notice that in the spirals tasks, FNO1D is initialized using more points ($5$ or $10$ intead of $2$ as in the Leray-Schauder neural operator). Therefore, FNO1D is given an advantage in initialization. For the Burgers’ equation, FNO2D was seen to perform significantly worse than ANIE in [ANIE], and we have therefore compared only with the latter in the present article for a more concise treatment.

### v3

@@ 3. Algorithm
In this section, we describe more in detail the concrete resulting algorithm from the previous theoretical results. Our deep learning algorithm is based on constructing a function input from the data available for initialization (e.g. an initial condition or boundary values, depending on the formulation of the problem), and (nonlinearly) project the input on a basis of functions modeled with neural networks. The projection maps are Leray-Schauder maps, defined through Equation (1). The maps $\mu_{i}$ are either fixed as in Equation (3) (for some choice of norm), or are learned as in Theorem 2.2 and Theorem 2.7. In fact, while our experiments show that the assignment in Equation (3) gives a functioning model, we have found that hyperparameter fine tuning is more complex in this setting, and simply learning the $\mu_{i}$ maps, which in turn means that we learn the Leray-Schauder maps as well, makes training much simpler.
Following the results in Section 2, the model consists of two basis sets of neural networks, and a neural network $f_{n,m}$ that operates on the spaces spanned by the bases. We point out some strong similarities with the DeepONet model, where one has a basis of neural networks (the trunk networks), and a set of functionals that project on such basis. In the present case, one can think of the first set of Leray-Schauder approximations and $f_{n,m}$ as playing the role of the projection functionals in DeepONet, and the second set of Leray-Schauder approximations as playing the role of the trunk neural networks. However, the resulting model does not seem to be equivalent to DeepONet, as it is not a priori clear whether the functionals of DeepONets can always decomposed into a Leray-Schauder approximation plus a feedforward neural network $f_{n,m}$.
Since we also learn the functions $\mu_{i}$ in Equation (3), in practice we consider convolutional neural networks $\mu_{i}:C(K,\mathbb{R}^{M})\longrightarrow\mathbb{R}^{+}_{0}$. We use CNNs to implement the $\mu_{i}$ neural networks because they mimic an integration procedure which can be performed in higher dimensions. They take a discretized input function and produce a numerical value (which can be constrained to be nonnegative by choice of the final nonlinearity to ensure increased stability).
We summarize the construction by means of the following commutative diagram
[…]
where $T^{\theta}_{n,m}$ represents the introduced model, which we will call the Leray-Schauder Neural Operator for brevity. The map $T_{n,m}$ is obtained by considering the restriction of $P_{m}T$ on $E_{n}$, and it is drawn dashed because $f_{n,m}$ only approximates this map on $\mathbb{K}\cap\bar{B}_{\rho}$ as above, and therefore the diagram would not strictly commute with such $T_{n,m}$. We have indicated the identifications $\phi_{n}$ and $\phi_{m}$ by equality symbols. The neural networks $\{h_{j}\}$ that span $E_{m}$ work in a fashion similar to the trunk neural networks of DeepONet, while the map $P_{n}$ is the approximation to the Leray-Schauder mapping $P_{n}$ of Section 2. The composition $\hat{\phi}_{m}^{-1}f_{n,m}\phi_{n}P_{n}$ in this diagram would correspond to the functionals used in DeepONet.
@@
which is the same operator used in the approximation of Theorem 2.7, where $F^{i}$ has been named $\mu^{i}$ to mimic the definition in (3). Since $P_{n}(\mathbf{y})$ can be evaluated through the networks $g_{i}$, it is clear that $P_{n}$ takes a function as input, obtained via initialization data, and produces a function as output. This function $P_{n}(\mathbf{y})$ lives inside the span $\langle g_{1},\ldots,g_{n}\rangle$. In fact, $P_{n}(\mathbf{y})$ is a linear combination of $g_{i}$, with coefficients $q_{i}:=\frac{\mu^{i}(\mathbf{y})}{\sum_{j=1}^{n}\mu^{j}(\mathbf{y})}$. The vector of coefficients $\mathbf{q}=(q_{1},\ldots,q_{n})^{T}\in\mathbb{R}^{n}$ is therefore the input of the network $f_{n,m}$, following the proof of Theorem 2.7.
We compute $\mathbf{b}=f(\mathbf{q})$, and $\mathbf{b}$ is used to take a linear combination of the neural networks $h_{i}$, which we set $\psi=\sum_{i=1}^{m}b_{i}h_{i}$. We can then evaluate $\psi(\mathbf{x},t)$ and compute the loss $\|\psi(\mathbf{x},t)-\mathbf{z}(\mathbf{x},t)\|$, where $\mathbf{z}$ is the target data which we are predicting.
This is summarized in Algorithm 1.
@@
1:Neural networks $\{g_{i}\}_{i=1}^{n}$, $\{h_{j}\}_{j=1}^{m}$ and $f_{n,m}$
2:Approximation $\phi_{m}^{-1}f_{n}\phi_{n}\tilde{P}_{n}$ of operator $T$
@@
6:Apply $f_{n,m}$ on $\mathbf{q}$: $f_{n,m}(\mathbf{q})=\mathbf{b}$
7:Output function, using $\phi^{-1}_{m}$, is $\psi(\mathbf{x},t)=\sum_{i=m}^{n}b_{i}h_{i}(\mathbf{x},t)$
@@ 4. Experiments
To test whether the model is stable under change of domain grid sampling, we consider two related situations. For the IE Spirals dataset, we consider the problem of interpolation. This means that the model is trained over a downsampled function, but tested on the full grid samples. It is seen that the model is stable under this upsampling procedure, due to the fact that the training procedure is effectively grid-independent. For the Burgers’ dataset, we consider two different grid sampling sizes for training. We see that the model accuracy is substantially independent of this sampling specifications. It was seen in [ANIE] that several models desplayed various behaviors under change of sampling size. While the ANIE model mitigated this via a Monte Carlo sampling procedure in the iterations of the solver, the present model is substantially more stable without the need of performing regularization procedures. In addition, the computational cost of the model is substantially unchanged when the grid size is varied. In fact, the main computational cost is related to the number of neural networks used for the two bases, and this value depends only on the problem at hand, and not on the grid size.
@@
Leray-Schauder $1\pm 0.0005 0.0011±0.00$$1\pm 0.0005$ $0017\pm 0.0012 0.0$$0017\pm 0.0011$
ANIE $0014\pm 0.0002 0.0$$015\pm 0.0003 0.0$$016\pm 0.0016 0.0$$0017\pm 0.0018$
Spectral NIE $0022\pm 0.0016 0.0$$0028\pm 0.0032$
FNO1D (init 5) $0292\pm 0.0285 0.0$$0994\pm 0.1207$
@@
$\mathbf{y}(t)=\int_{0}^{t}\begin{bmatrix}\cos 2\pi(t-s)&-\sin 2\pi(t-s)\\ -\sin 2\pi(t-s)&-\cos 2\pi(t-s)\end{bmatrix}\tanh(2\pi\mathbf{y}(s))dx+\mathbf{z}_{0}+\begin{bmatrix}\cos(t)\\ \cos(t+\pi)\end{bmatrix}$,
@@
In the experiments, we find that the model is comparable with the state of the art models (as in the experiments in [ANIE]) on both datasets. For the IE spirals, we have included two experiments, one on prediction of the original dynamics, and another one on an interpolation task where the model has access to half the points during training and predicts all of them during evaluation. For the Burgers’ dataset we have experimented on two spatial resolutions $s=256$ and $s=512$. An example of ground truth dynamics for the Burgers’ equation is found in Figure 1, while the corresponding model’s prediction is in Figure 2. The vertical direction indicates space and the horizontal direction indicates time.
@@
We notice that in the spirals tasks, FNO1D is initialized using more points ($5$ or $10$ intead of $2$ as in the Leray-Schauder neural operator). Therefore, FNO1D is given more points for initialization. For the Burgers’ equation, FNO2D was seen to perform significantly worse than ANIE in [ANIE], and we have therefore compared only with the latter in the present article for a more concise treatment.